elena is 6 years younger than her brother.the sum of their ages is 30 how old is elena oh you have to guess and test
Answer by wilft1(217)

guess and test all you want, the formula for this would be
x + x - 6 = 30
add 6 to both sides
x + x = 36
combine like terms
2x = 36
divide by 2
x = 18
x - 6 = 12
elena is 12 and her brother is 18